UWANTEDSKILL:

There are 60 half dollars and nickels. If they total $12 how many of each are there?

curriful:

does this mean there is 60 EACH or just 60 in all?

curriful:

bc if it was together, there should be 18 half dollars and 60 nickels

curriful:

for each one, there should be 24 half dollars and 400 nickels

curriful:

wait now it seems obvious which one it is..

curriful:

yea, its 18 half dollars and 60 nickels

YourlocalRandom2000:

5x + 5y = 300 (equation 4) Subtracting equation 4 from equation 3, we get: 45x = 900 Dividing both sides by 45, we get: x = 20 Substituting x = 20 into equation 1, we get: 20 + y = 60 Subtracting 20 from both sides, we get: y = 40 there are 20 half dollars and 40 nickels.

OLIVER69:

I think y'all are making this too complicated. There are 60 total of both half dollars and nickles, if their total is 12 all together, then you simply just do 60 divide by 12, and your answer would be 5.

surjithayer:

let number of half dollars=x number of nickels=y x+y=60 \[\frac{ 1 }{ 2 }x+\frac{ 5 }{ 100 }y=12\] \[\frac{ 1 }{ 2 }x+\frac{ 1 }{ 20 }y=12\] multiply by 20 \[10x+y=12 \times12\] 10x+(60-x)=240 9x=240-60 9x=180 x=180/9=20 y=60-20=40 number of half dollars=20 number of nickels=40